OpenStudy (anonymous):

Solve the logarithmic equation log(5+x)-log(x-2)=log2

OpenStudy (anonymous):

\[\log _{a}x-\log _{a}y=\log _{a}x/y\] applying that here... log((5+x)/(x-2))=log 2 since log is being taken of both sides you can just equate the two (5+x)/(x-2)=2 know what to do from here?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Multiply sides by x-2 5+x=2(x-2) 5+x=2x-4 9+x=2x 9=x